# Quark-versus-gluon tagging in CMS Open Data with CWoLa and TopicFlow

### v1

@@ Abstract
We use the CMS Open Data to examine the performance of weakly-supervised learning for tagging quark and gluon jets at the LHC. We target $Z$+jet and dijet events as respective quark- and gluon-enriched mixtures and derive samples both from data taken in 2011 at 7 TeV, and from Monte Carlo. CWoLa and TopicFlow models are trained on real data and compared to fully-supervised classifiers trained on simulation. In order to obtain estimates for the discrimination power in real data, we consider three different estimates of the quark/gluon mixture fractions in the data. Compared to when the models are evaluated on simulation, we find reversed rankings for the fully- and weakly-supervised approaches. Further, these rankings based on data are robust to the estimate of the mixture fraction in the test set. Finally, we use TopicFlow to smooth statistical fluctuations in the small testing set, and to provide uncertainty on the performance in real data.
@@ Introduction
Discriminating between the quark or gluon origins of hadronic jets is a classic and difficult problem in the phenomenology of quantum chromodynamics (QCD) at particle colliders, which has been extensively studied experimentally [1, 2, 3, 4, 5, 6, 7, 8, 9]. There are multiple reasons it would be desirable to do this. From the perspective of Beyond the Standard Model physics, tagging hadronic decays of potential resonances could be helpful in reconstructing the underlying UV theory. From a Standard Model (SM) perspective, studies that aim to isolate Higgs production from vector boson fusion, where a Higgs is produced in association with two quark jets, must determine the contamination from gluon fusion, which can involve Higgs production with multiple gluon jets.
@@
Like other jet classification tasks, there has been extensive study of neural-network-based classifiers applied to quark/gluon discrimination [10, 6, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. These models, built on modern deep-learning architectures, are able to achieve discrimination power far exceeding traditional approaches based on jet substructure observables. However, typically they are trained and tested in a fully-supervised setting, making use of parton-level information in Monte-Carlo (MC). Given that these taggers owe their strength to a sensitivity to the detailed correlations in the substructure of a jet, they are at risk of overfitting to mismodelled features in Monte Carlo (MC) samples. In particular, quark/gluon discrimination is known to be highly sensitive to non-perturbative effects in QCD [24, 25]. Combining this with the fact that contemporary simulations for parton showering and hadronisation have large theoretical uncertainties [26], we are motivated to explore methods for training and evaluating quark/gluon taggers on real data.
@@
Such methods can be realised within the paradigm of weakly-supervised learning, which broadly covers methods that take as input datasets with noisy labels or, equivalently, mixed classes. Such datasets are generally easier to obtain than the completely separated samples required for full supervision. Indeed, with motivated phase-space selections, one can isolate regions enriched with quark or gluon jets [27] at the parton level. These selections can then be applied in data, yielding input samples for weakly-supervised algorithms. One such method is Classification Without Labels (CWoLa, pronounced “koala”) [28, 29], where a classifier is tasked with discriminating two datasets of different mixture proportions. Under certain assumptions, this kind of classifier can be expected to match the performance of a fully-supervised model. In contrast to previous weakly-supervised paradigms [30], this is possible even without knowledge of the class proportions. Previous works employing CWoLa for quark/gluon classification include Refs. [16, 31]. [^1] See also Ref. [32] for an unsupervised approach.
@@
More recent work on jet topics [33, 34, 31] offers a framework within which the mixture fractions can be estimated in data, leading to predictions for the individual quark- and gluon-like distributions. This technique provides an ‘operational’ definition of the terms ‘quark jet’ and ‘gluon jet’ and has seen application in both experimental [35, 36, 37, 38] and theoretical [39, 40, 41] settings. In a previous work [42] some of the authors introduced a deep generative approach to jet topics, called TopicFlow. There, we showed that normalising flows can demix quark and gluon distributions in many dimensions and demonstrated their advantage by smoothing statistics with oversampling. While TopicFlow is compatible with use on real data, previously it was only tested on simulated collision events at the Large Hadron Collider (LHC).
@@
The CMS collaboration has made a wide range of data collected by the LHC publicly available through the CERN Open Data Portal [43]. This now includes fully reconstructed collision data as well as detector-simulated MC samples, paving the way for exploratory analysis strategies, measurements, and new data formats and frameworks to be investigated outside of the experimental collaborations. Work in these directions has already begun, with jet studies [44, 45, 31, 46, 47] as well as machine-learning studies [48, 49, 50, 51] and new search strategies and techniques [52, 53, 54, 55, 56, 57].
@@
In this paper, we make use of the CMS Open Data to extend investigations of weakly-supervised learning using TopicFlow. A previous work [31] has performed a similar study, defining quark- and gluon-enriched samples by partitioning QCD jets in pseudorapidity (see also  [9]). Another possibility, recently advocated in [58], is to combine measurements made at different energies. In contrast to those works, and similarly to CMS, we use distinct event topologies for our sources of enriched samples. Specifically, we isolate $Z+$jet and dijet samples which are respectively expected to be enriched in quark jets and gluon jets. Using both simulated and real events, we extract jet topics and train CWoLa and TopicFlow models. Significance improvement curves evaluated on MC are compared with estimates for the curves in data, revealing reversed rankings for the fully- and weakly-supervised approaches.
@@
The remainder of the paper is organised as follows. In Section I we detail the selections that define our enriched samples in simulation and collision data. Section II describes the weakly-supervised methods that we apply to the datasets, with specific implementations outlined in Section III. We present the results of our study in Section IV before providing concluding remarks in Section V.
I Datasets and Event Selection
@@
The experimental data we use for our analysis was collected by the CMS experiment in 2011 and made publicly available through the CERN Open Data Portal [43] in 2016. It includes 2.5 $2.5\text{ fb}^{-1}$ of $7$ TeV proton–proton collisions collected in Run 2011A, as well as simulated events which have been passed through the CMS detector simulation. This was the second of a number of batches of data published regularly by the collaboration since 2014. Currently, more than 2 PB from the 2010–2012 run period have been made publicly available, including all data collected in 2010 and 2011.
Figure 1: Validation of the selection cuts applied to the $Z+$jet samples. The panels show a) leading jet $p_{T}$, b) dimuon $p_{T}$, c) dimuon invariant mass and d) leading muon-jet separation $\Delta R$. The DY Monte Carlo (from Pythia) is in purple, the data is the black points. The error bars show statistical uncertainties. All distributions are normalised to unit area.
Figure 2: Validation of the selection cuts applied to the dijet samples. The left-hand plot shows the leading jet $p_{T}$ and right-hand one the average dijet $p_{T}$. The QCD Monte Carlo (from Pythia) is in green and the data is the black points. The error bars show statistical uncertainties. All distributions are normalised by area.
@@
Our study makes use of four datasets: quark-enriched and gluon-enriched samples from both collision data and Monte Carlo (MC) simulation. For the gluon-enriched datasets, we use the Jet primary dataset from RunA of 2011 [59] and the SM-exclusive QCD simulation generated with $Z2$-tuned Pythia 6 [60]. These simulated data are accessed through separate samples organised by $p_{T}$ ranges [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. All of the jet triggers in the MC fire with unit prescales. We consider jets with $p_{T}>30$ GeV and $|\eta|<2.1$ that satisfy medium jet quality criteria (JQC) [74]. Events with fewer than two such jets are discarded and we also require the DiJetAve30 high-level trigger to fire in the collision events. For the quark-enriched datasets, we use the DoubleMu primary dataset from RunA and RunB of 2011 [75] and the SM-inclusive Drell–Yan (DY) MC, also generated with $Z2$-tuned Pythia 6 [76]. We take events that contain two global tracker muons [77] satisfying $p_{T,\mu}>26$ GeV and $|\eta_{\mu}|<2.4$. The leading dimuon system is identified as a $Z$ candidate and required to have 70 GeV $\leq m_{\mu\mu}\leq$ 110 GeV. After discarding jets that overlap with either muon within their radii, we use the same jet definition and cuts as for the gluon-enriched sample. Events are further subjected to $\Delta\phi(j_{0},\mu\mu)>2$ and $\frac{|p_{T,\mu\mu}-p_{T,j_{0}}|}{p_{T,\mu\mu}+p_{T,j_{0}}}<0.3$ where $j_{0}$ refers to the leading jet [8]. For the DoubleMu dataset, we require a firing Mu13_Mu8 high-level trigger, which is not prescaled. In all datasets, we select events where the leading jet has $p_{T}$ in the range $[50,80]$ GeV. In the two simulated datasets, we remove jets that are not assigned a parton match, which is the case for approximately 15% of the jets in QCD MC and 3% in DY MC. Finally, the parton labels are used to construct ‘pure’ quark and gluon datasets from the QCD simulation which will facilitate supervised training. The number of events passing selection in each dataset is given in Table 1.
@@
Table 1: Summary of the number of events remaining in each dataset as well as the flavour of the leading jet (MC only) after all selections described in Section I.
In Figure 1 we show validation plots for our event selection and workflow for the $Z+$jet samples. We compare the DY Monte Carlo (in purple) with the data from the DoubleMu Primary dataset (black data-points). All distributions have been normalised by area. Clockwise from top-left we show the leading jet $p_{T}$, the dimuon $p_{T}$, the separation $\Delta R$ between the leading muon and jet, and the dimuon invariant mass.
We find excellent agreement in the leading jet $p_{T}$ and dimuon $p_{T}$, although statistics start to become an issue at higher muon $p_{T}$ values. The dimuon invariant mass is peaked at $m_{Z}$, although the shape is not perfectly modelled. We also observe some small discrepancies at large muon-jet separation. We checked that these variables are not correlated to the substructure observables of interest to our study.
Figure 2 shows two distributions derived from the QCD Monte Carlo and the Jet Primary Dataset. The left-hand plot shows the leading jet $p_{T}$ and right-hand one the average dijet $p_{T}$. Again, we observe good agreement between the simulation and data. Studies of jet physics using the CMS Open Data have also been undertaken in a number of papers by the MIT group. We have compared our processed data with that made publicly available through Ref. [54] in Appendix B.
II Weakly-supervised methods
II.1 Classification without labels
The classification without labels (CWoLa) paradigm was first explored for high energy physics in [28]. In this picture, we assume that two mixtures $M_{1}$ and $M_{2}$ are differently proportioned combinations of the same underlying ‘quark’ and ‘gluon’ components such that we can write distributions for any observable $x$ as
@@ II.1 Classification without labels
In practice, there are various reasons for which a weakly-supervised CWoLa classifier may not be able to match the performance of a fully-supervised classifier. Firstly, the above arguments only guarantee that the two classifiers share an optimum. Since neural-network training can only ever approximate the optimal classifier, limitations such as the available training statistics can have impact. Additionally, the framework relies on the assumption that the two mixtures contain the same $p_{Q}$ and $p_{G}$, called sample independence. Depending on the source of each mixture, there may be some statistical difference between quarks in $M_{1}$ and those in $M_{2}$, for example. It is known that soft colour correlations can cause sample dependence for $Z+$jet and dijet samples, although previous studies have observed this to be a small effect, particularly for small-radius jets [78]. Detector effects may also introduce large differences. [^2] Detector-induced sample dependence may be alleviated through unfolding methods, as was applied in [31]. We will assess the level to which sample independence is a valid assumption for our datasets in Section IV.
II.2 Jet topics

### v2

@@ Abstract
We use the CMS Open Data to examine the performance of weakly-supervised learning for tagging quark and gluon jets at the LHC. We target ZZ+jet and dijet events as respective quark- and gluon-enriched mixtures and derive samples both from data taken in 2011 at 7 TeV, and from Monte Carlo. CWoLa and TopicFlow models are trained on real data and compared to fully-supervised classifiers trained on simulation. In order to obtain estimates for the discrimination power in real data, we consider three different estimates of the quark/gluon mixture fractions in the data. Compared to when the models are evaluated on simulation, we find reversed rankings for the fully- and weakly-supervised approaches. Further, these rankings based on data are robust to the estimate of the mixture fraction in the test set. Finally, we use TopicFlow to smooth statistical fluctuations in the small testing set, and to provide uncertainty on the performance in real data.
@@ Introduction
Discriminating between the quark or gluon origins of hadronic jets is a classic and difficult problem in the phenomenology of quantum chromodynamics (QCD) at particle colliders, which has been extensively studied experimentally DELPHI:1999gah ; OPAL:1999jkz ; CDF:2005prv ; ATLAS:2016wzt ; ATLAS:2017nma ; ATLAS:2017dfg ; CMS:2017yer ; CMS:2021iwu ; ATLAS:2023pdx . There are multiple reasons it would be desirable to do this. From the perspective of Beyond the Standard Model physics, tagging hadronic decays of potential resonances could be helpful in reconstructing the underlying UV theory. From a Standard Model (SM) perspective, studies that aim to isolate Higgs production from vector boson fusion, where a Higgs is produced in association with two quark jets, must determine the contamination from gluon fusion, which can involve Higgs production with multiple gluon jets.
@@
Like other jet classification tasks, there has been extensive study of neural-network-based classifiers applied to quark/gluon discrimination Komiske:2016rsd ; ATLAS:2017dfg ; Cheng:2017rdo ; Luo:2017ncs ; Kasieczka:2018lwf ; Komiske:2018cqr ; Lee:2019cad ; Lee:2019ssx ; Moreno:2019bmu ; Qu:2019gqs ; Mikuni:2020wpr ; Dreyer:2020brq ; Bogatskiy:2022czk ; Butter:2022xyj ; Qu:2022mxj ; He:2023cfc . These models, built on modern deep-learning architectures, are able to achieve discrimination power far exceeding traditional approaches based on jet substructure observables. However, typically they are trained and tested in a fully-supervised setting, making use of parton-level information in Monte-Carlo (MC). Given that these taggers owe their strength to a sensitivity to the detailed correlations in the substructure of a jet, they are at risk of overfitting to mismodelled features in Monte Carlo (MC) samples. In particular, quark/gluon discrimination is known to be highly sensitive to non-perturbative effects in QCD Gras:2017jty ; Dreyer:2021hhr . Combining this with the fact that contemporary simulations for parton showering and hadronisation have large theoretical uncertainties mo:2017gzp , we are motivated to explore methods for training and evaluating quark/gluon taggers on real data.
@@
Such methods can be realised within the paradigm of weakly-supervised learning, which broadly covers methods that take as input datasets with noisy labels or, equivalently, mixed classes. Such datasets are generally easier to obtain than the completely separated samples required for full supervision. Indeed, with motivated phase-space selections, one can isolate regions enriched with quark or gluon jets Gallicchio:2011xc at the parton level. These selections can then be applied in data, yielding input samples for weakly-supervised algorithms. One such method is Classification Without Labels (CWoLa, pronounced “koala”) Metodiev:2017vrx ; Komiske:2018oaa , where a classifier is tasked with discriminating two datasets of different mixture proportions. Under certain assumptions, this kind of classifier can be expected to match the performance of a fully-supervised model. In contrast to previous weakly-supervised paradigms Dery:2017fap , this is possible even without knowledge of the class proportions. Previous works employing CWoLa for quark/gluon classification include Refs. Lee:2019ssx ; ATLAS:2020iwa ; Komiske:2022vxg ; CMS:2024nsz . [^1] See also Ref. Alvarez:2021zje for an unsupervised approach.
@@
More recent work on jet topics Metodiev:2018ftz ; Komiske:2018vkc ; Komiske:2022vxg offers a framework within which the mixture fractions can be estimated in data, leading to predictions for the individual quark- and gluon-like distributions. This technique provides an ‘operational’ definition of the terms ‘quark jet’ and ‘gluon jet’ and has seen application in both experimental ATLAS:2019rqw ; Brewer:2020och ; Ying:2022jvy ; LeBlanc:2022bwd and theoretical Larkoski:2019nwj ; Stewart:2022ari ; Takacs:2021bpv settings. In a previous work Dolan:2022ikg some of the authors introduced a deep generative approach to jet topics, called TopicFlow. There, we showed that normalising flows can demix quark and gluon distributions in many dimensions and demonstrated their advantage by smoothing statistics with oversampling. While TopicFlow is compatible with use on real data, previously it was only tested on simulated collision events at the Large Hadron Collider (LHC).
@@
The CMS collaboration has made a wide range of data collected by the LHC publicly available through the CERN Open Data Portal cernopendataportal . This now includes fully-reconstructed collision data as well as detector-simulated MC samples, paving the way for exploratory analysis strategies, measurements, and new data formats and frameworks to be investigated outside of the experimental collaborations. Work in these directions has already begun, with jet studies Larkoski:2017bvj ; Tripathee:2017ybi ; Komiske:2022vxg ; Komiske:2022enw ; Arias:2023wai as well as machine-learning studies Madrazo:2017qgh ; Andrews:2018nwy ; Andrews:2019faz ; knapp2020adversarially and new search strategies and techniques PaktinatMehdiabadi:2019ujl ; Cesarotti:2019nax ; Komiske:2019jim ; Lester:2019bso ; Apyan:2019ybx ; An:2021yqd .
@@
In this paper, we make use of the CMS Open Data to extend investigations of weakly-supervised learning using TopicFlow. A previous work Komiske:2022vxg has performed a similar study, defining quark- and gluon-enriched samples by partitioning QCD jets in pseudorapidity (see also  ATLAS:2023pdx ). Another possibility, recently advocated in Baron:2023hkp , is to combine measurements made at different energies. In contrast to those works, and similarly to CMS, we use distinct event topologies for our sources of enriched samples. Specifically, we isolate Z+Z+jet and dijet samples which are respectively expected to be enriched in quark jets and gluon jets. Using both simulated and real events, we extract jet topics and train CWoLa and TopicFlow models. Significance improvement curves evaluated on MC are compared with estimates for the curves in data, revealing reversed rankings for the fully- and weakly-supervised approaches.
@@
The remainder of the paper is organised as follows. In Section 1 we detail the selections that define our enriched samples in simulation and collision data. Section 2 describes the weakly-supervised methods that we apply to the datasets, with specific implementations outlined in Section 3. We present the results of our study in Section 4 before providing concluding remarks in Section 5.
1 Datasets and Event Selection
@@
The experimental data we use for our analysis was collected by the CMS experiment in 2011 and made publicly available through the CERN Open Data Portal cernopendataportal in 2016. It includes $2.5\text{ fb}^{-1}$ of $7$ TeV proton–proton collisions collected in Run 2011A, as well as simulated events which have been passed through the CMS detector simulation. This was the second of a number of batches of data published regularly by the collaboration since 2014. Currently, more than 2 PB from the 2010–2012 run period have been made publicly available, including all data collected in 2010 and 2011.
@@
Our study makes use of four datasets: quark-enriched and gluon-enriched samples from both collision data and Monte Carlo (MC) simulation. For the gluon-enriched datasets, we use the Jet primary dataset from RunA of 2011 jet2011A and the SM-exclusive QCD simulation generated with $Z2$-tuned Pythia 6 Sjostrand:2006za . These simulated data are accessed through separate samples organised by $p_{T}$ ranges qcdsim15to30 ; qcdsim30to50 ; qcdsim50to80 ; qcdsim80to120 ; qcdsim120to170 ; qcdsim170to300 ; qcdsim300to470 ; qcdsim470to600 ; qcdsim600to800 ; qcdsim800to1000 ; qcdsim1000to1400 ; qcdsim1400to1800 ; qcdsim1800 . All of the jet triggers in the MC fire with unit prescales. A detailed discussion of the CMS Run 1 Trigger system can be found in CMS:2016ngn . We consider jets with $p_{T}>30$ GeV and $|\eta|<2.1$ that satisfy medium jet quality criteria (JQC) CMS:2011shu . Events with fewer than two such jets are discarded and we also require the DiJetAve30 high-level trigger to fire in the collision events. For the quark-enriched datasets, we use the DoubleMu primary dataset from RunA and RunB of 2011 doublemu2011A and the SM-inclusive Drell–Yan (DY) MC, also generated with $Z2$-tuned Pythia 6 dymc . We take events that contain two global tracker muons CMS:2018rym satisfying $p_{T,\mu}>26$ GeV and $|\eta_{\mu}|<2.4$. The leading dimuon system is identified as a $Z$ candidate and required to have 70 GeV $\leq m_{\mu\mu}\leq$ 110 GeV. After discarding jets that overlap with either muon within their radii, we use the same jet definition and cuts as for the gluon-enriched sample. Events are further subjected to $\Delta\phi(j_{0},\mu\mu)>2$, ensuring that the selected events are dominantly two-body processes ($Z+j$) with minimal additional jet activity. We also apply a $p_{T}$ asymmetry cut $\frac{|p_{T,\mu\mu}-p_{T,j_{0}}|}{p_{T,\mu\mu}+p_{T,j_{0}}}<0.3$ where $j_{0}$ refers to the leading jet CMS:2021iwu . This limits contributions from large QCD corrections associated with events where $p_{j1,T}\gg p_{Z,T}$  Rubin:2010xp . For the DoubleMu dataset, we require a firing Mu13_Mu8 high-level trigger, which is not prescaled. In all datasets, we select events where the leading jet has $p_{T}$ in the range $[50,80]$ GeV. In the two simulated datasets, we remove jets that are not assigned a parton match, which is the case for approximately 15% of the jets in QCD MC and 3% in DY MC (see Appendix C for further discussion). Finally, the parton labels are used to construct ‘pure’ quark and gluon datasets from the QCD simulation which will facilitate supervised training. The number of events passing selection in each dataset is given in Table 1.
@@
Table 1: Summary of the number of events remaining in each dataset as well as the flavour of the leading jet (MC only) after all selections described in Section 1.
Figure 1: Validation of the selection cuts applied to the Z+Z+jet samples. The panels show a) leading jet $p_{T}$, b) dimuon $p_{T}$, c) dimuon invariant mass and d) leading muon-jet separation $\Delta R$. The DY Monte Carlo (from Pythia) is in purple, the data is the black points. The error bars show statistical uncertainties. All distributions are normalised to unit area.
In Figure 1 we show validation plots for our event selection and workflow for the Z+Z+jet samples. We compare the DY Monte Carlo (in purple) with the data from the DoubleMu Primary dataset (black data-points). All distributions have been normalised by area. Clockwise from top-left we show the leading jet $p_{T}$, the dimuon $p_{T}$, the separation $\Delta R$ between the leading muon and jet, and the dimuon invariant mass. Here, and throughout this paper, uncertainty bands are statistics-only. We find excellent agreement in the leading jet $p_{T}$ and dimuon $p_{T}$, although statistics start to become an issue at higher muon $p_{T}$ values. The dimuon invariant mass is peaked at $m_{Z}$, although the shape is not perfectly modelled. We also observe some small discrepancies at large muon-jet separation.
Figure 2: Validation of the selection cuts applied to the dijet samples for LABEL:sub@fig:dijet_jetPt.pdf the leading jet $p_{T}$ and LABEL:sub@fig:dijet_aveDijetPt.pdf the average dijet $p_{T}$. The QCD Monte Carlo (from Pythia) is in green and the data is the black points. The error bars show statistical uncertainties. All distributions are normalised by area.
Figure 2 shows two distributions derived from the QCD Monte Carlo and the Jet Primary Dataset. The left-hand plot shows the leading jet $p_{T}$ and right-hand one the average dijet $p_{T}$. Again, we observe good agreement between the simulation and data.
For our study, we also consider jet substructure observables used for quark/gluon discrimination in previous CMS works CMS-PAS-JME-13-002 ; CMS-PAS-JME-16-003 . We consider the multiplicity of the jet constituents, the jet mass, the jet fragmentation distribution $p_{T}^{D}$ and the major axis of the jet $\sigma_{1}$. Definitions for the latter two observables can be found in Appendix B. In Figure 3, we compare the data and MC distributions for the leading jet substructure. The left column shows dijet events, which exhibit discrepancies on the order of 25%, notably for multiplicity and $p_{T}^{D}$. The agreement is somewhat better for Z+Z+jet events in the right column, though systematic deviations remain. Given that we only have access to statistical uncertainties in MC, we cannot determine whether the observed biases are covered by varying systematics such as the parton shower or renormalisation/factorisation scales. We checked that these substructure variables are not correlated to the event-level observables that also displayed biases in Figure 1.
Figure 3: Data/MC comparison for jet substructure observables in dijet events (LABEL:sub@fig:dijet_nConstits, LABEL:sub@fig:dijet_sigma_1, LABEL:sub@fig:dijet_ptD) and Z+Z+jet events (LABEL:sub@fig:zjet_nConstits, LABEL:sub@fig:zjet_sigma_1, LABEL:sub@fig:zjet_ptD).
Studies of jet physics using the CMS Open Data have also been undertaken in a number of papers by the MIT group. We have compared our processed data with that made publicly available through Ref. Komiske:2019jim in Appendix D.
2 Weakly-supervised methods
2.1 Classification without labels
The classification without labels (CWoLa) paradigm was first explored for high energy physics in Metodiev:2017vrx . In this picture, we assume that two mixtures $M_{1}$ and $M_{2}$ are differently proportioned combinations of the same underlying ‘quark’ and ‘gluon’ components such that we can write distributions for any observable $x$ as
@@ 2.1 Classification without labels
In practice, there are various reasons for which a weakly-supervised CWoLa classifier may not be able to match the performance of a fully-supervised classifier. Firstly, the above arguments only guarantee that the two classifiers share an optimum. Since neural-network training can only ever approximate the optimal classifier, limitations such as the available training statistics can have impact. Additionally, the framework relies on the assumption that the two mixtures contain the same $p_{Q}$ and $p_{G}$, called sample independence. Depending on the source of each mixture, there may be some statistical difference between quarks in $M_{1}$ and those in $M_{2}$, for example. It is known that soft colour correlations can cause sample dependence for Z+Z+jet and dijet samples, although previous studies have observed this to be a small effect, particularly for small-radius jets Bright-Thonney:2018mxq . Detector effects may also introduce large differences. [^2] Detector-induced sample dependence may be alleviated through unfolding methods, as was applied in Komiske:2022vxg . We will assess the level to which sample independence is a valid assumption for our datasets in Section 4.
2.2 Jet topics